OpenStudy (anonymous):

The functions f(x) and g(x) are described below: f(x) = 7x + 9 g(x) = 7x - 4 The graph of g(x) is obtained by shifting down the graph of f(x) by _____ units. I got 5 as the answer, am I correct?

OpenStudy (anonymous):

No, it's 13 units . Y=mx+b You start at b which for f(x) would be 9 for g(x) it's -4 9-(-4)=9+4=13

OpenStudy (anonymous):

your right @Xomoneek